OpenStudy (anonymous):

Dulce adds 1 1/3 cup oil to a recipe and then realize that should have added 1 1/4 cup oil. about how much more oil does she need to add? Explain

OpenStudy (anonymous):

Hi christyx, If Dulce adds 1 1/3 cups of oil to a recipe and finds that they need 1 1/4, it would be the case that they have added too much, not too little. To see this, note that\[1\frac{1}{3}-1\frac{1}{4}=\frac{4}{3}-\frac{5}{4}=\frac{4 \times 4 - 3 \times 5}{3 \times 4}=\frac{1}{12}\]which is greater than 0. That means the amount in the recipe is greater than what's needed. So Dulce shouldn't add any more...Dulce should consider starting again, or adjusting the amounts of other ingredients :)